OpenStudy (anonymous):

Write an equation for the line that passes through (-24,-11) and (-8,-3).

OpenStudy (anonymous):

find the slope of the line using \(\huge{m={{y_2-y_1}\over{x_2-x_1}}}\). once you've found the slope plug in the slope, and one of the two given set of points into the point-slope formula \(\huge y-y_1=m(x-x_1)\).

OpenStudy (anonymous):

So, far I got m=2

OpenStudy (anonymous):

What do I do next?

OpenStudy (anonymous):

try doing the slope again:\[m={{(-3)-(-11)}\over{(-8)-(-24)}}\]\[m={{-3+11}\over{-8+24}}\]\[m={{8}\over{16}}\]\[m={1\over2}\]

OpenStudy (anonymous):

remember, its the difference in \(y\) over the difference in \(x\) :)

OpenStudy (anonymous):

now that you have that, just plug in one of the points along with the slope in the formula i gave above:\[y-(-11)={1\over2}(x-(-24))\]\[y+11={1\over2}(x+24)\]\[y+11={1\over2}x+12\]\[y={1\over2}x+1\]

OpenStudy (anonymous):

hope that helps! :) @smithjessica27

OpenStudy (anonymous):

yup, thanks :)